# The polytope of all $q$-rank functions

### v1

@@ 2. Preliminaries
The first results on $q$-polymatroids and their relation to coding theory can be read in [16] and [24]. The following definition is from [16, Def.4.1].
@@
The definition of $q$-polymatroid appeared in different forms in the literature. In [11, Def. 2.1] the same definition occurs with the only difference that the rank function takes only rational values. In this case we talk about rational $q$-polymatroids. Fundamental concepts regarding rational $q$-polymatroids have been investigated in [12] and [11]. A number $\mu\in\mathbb{Q}_{\geq 0}$ is called a denominator of $\rho$ and $\mathcal{M}$ if $\mu\rho(X)\in\mathbb{N}_{0}$ for all $X\in\mathcal{L}(E)$. The smallest denominator is called the principal denominator. A $q$-polymatroid with principal denominator $1$ (i.e., $\rho(X)\in\mathbb{N}_{0}$ for all $X$) is a $q$-matroid.
Next, for any $r\in\mathbb{N}$ a $(q,r)$-polymatroid as in [24, Def. 2] by Shiromoto, [9, Def. 1] and [3, Def. 1] can be turned into a rational $q$-polymatroid with denominator $r$ by dividing the rank function by $r$. Conversely, given a rational $q$-polymatroid with denominator $\mu$, then $(\mathcal{L}(E),\mu\rho)$ is a $(q,\lceil\mu\rceil)$-polymatroid. Finally, in [1, Def. 2.12] and [4, Def. 2.2] an $\mathcal{L}$-polymatroid is defined as a direct extension of the definition of the previous $q$-polymatroid.
We will use the following definition of independent spaces derived in [11, Def. 3.1].
@@
In [11, Def. 3.8], the authors define the $\mu$-bases of a space $V\in\mathcal{L}(E)$ as the inclusion-maximal $\mu$-independent subspaces of $V$. Furthermore, they show that all the $\mu$-bases of $V$ have the same rank value, which is equal to the rank of $V$.
@@
The following definition comes from [4, Def. 4.2].
@@ 4.2. The cyclic spaces
By following [14, Def. 7.7], we introduce the notion of full $q$-polymatroid.
@@ 4.3. The $\mu$-independent spaces
Remark 4.12.
By definition of $\mathcal{M}$ we clearly have that the $\mu=\mathrm{denom}(\lambda)$ always is a denominator for $\mathcal{M}$, but in general such a $\mu$ might not be principal. Nevertheless, Proposition 4.13 shows that there exist infinitely many cases in which the denominator of $\lambda$ is in fact a principal denominator for $\mathcal{M}$.
Proposition 4.13.
If $\mu=\text{denom}(\lambda)$ is prime, then it is principal.
@@
Let $\lambda=\frac{a}{b}$, where $b$ is a prime number and $\rho$ be the $q$-rank function of $\mathcal{M}$. By Theorem 3.8 we know that $\mathcal{M}$ corresponds to a rational point in $\mathcal{P}_{q}^{n}$, i.e., there exists a $X\in\mathcal{L}(E)$ such that
$\rho(X)=\lambda\rho_{1}(X)+(1-\lambda)\rho_{2}(X)=\frac{a\rho_{1}(X)+(b-a)\rho_{2}(X)}{b}=\frac{Z}{b}$
and $b$ does not divide $Z$. Moreover $\frac{Z}{b}$ cannot be simplified, since $b$ is a prime number. Let $\mu$ be the principal denominator of $\mathcal{M}$, meaning that $\mu\frac{Z}{b}$ is an integer. So, $b$ has to divide $\mu Z$. Since $b$ is prime, and it does not divide $Z$, it must divide $\mu$. Finally, since $\mu$ is principal and every value of $\rho$ can be written as rational number with denominator $b$, we get that $\mu=b$. ∎
@@ 5. The convex combination of paving q-matroids
[12, Prop 4.6.] Let $11\leq k\leq n-$ 1. In addition, let $\mathcal{S}\subseteq\mathcal{L}(E)_{k}$, such that $\dim(V\cap W)\leq k-2$, for every two distinct $V,W\in\mathcal{S}$. Define the map
@@
Therefore, if a $q$-polymatroid is paving according to Definition 5.4, then the dimension of every $\mu$-circuit is greater or equal than the rank of the $q$-polymatroid. A detailed discussion about $\mu$-independent spaces can be found in [11, Sec. 3-4].
@@ 7. The characteristic Puiseux polynomial
In [3, Sec. 3], the characteristic polynomial of an integer $(q,r)$-polymatroid was introduced and studied. Moreover the special case given by a $(q,1)$-polymatroid, i.e. a $q$-matroid, was then investigated in more detail in [19, Sec. 5]. In this section, we introduce a more general notion for rational $q$-polymatroids, inspired by [3]. However, this is not a polynomial anymore, but rather a truncated Puiseux series, i.e., a finite linear combination of monomials whose exponents may be rational and possibly negative. Afterwards, we discuss some properties in the context of $q$-polymatroids arising from convex combination of $q$-matroids.
@@
We want to emphasis that Theorem 7.4 does not say only that $\chi_{\mathcal{M}}$ is determined by either $\chi_{\mathcal{M}_{1}}$ or $\chi_{\mathcal{M}_{2}}$. By using [19, Thm. 5.8], we also observe that $\chi_{\mathcal{M}}$ can be computed by just using the flats of either $\mathcal{M}_{1}$ or $\mathcal{M}_{2}$. In other words, in order to determine $\chi_{\mathcal{M}}$ we do not need any concepts associated to the $q$-polymatroid $\mathcal{M}$, but rather only the well understood notion of flats of either of the $q$-matroids $\mathcal{M}_{1}$ or $\mathcal{M}_{2}$.
@@ 8. Representable q-polymatroid
Then $(\mathcal{L}(E),\rho)$ is a $q$-polymatroid [16, Theorem 5.3] and we denote it by $\mathcal{M}[C]$.
@@
Let $\rho:\mathcal{L}(E)\longrightarrow\mathbb{Z}_{\geq 0}$ be defined by $\rho(W):=k-\dim_{\mathbb{F}_{q^{m}}}(C(W)).$ Then $(\mathcal{L}(E),\rho)$ is a $q$-matroid [20, Thm. 24] and we also denote it by $\mathcal{M}[C]$. Note that in this case, $k$ is the dimension of $C$ over $\mathbb{F}_{q^{m}}$.
@@
[16, Prop. 6.2] Let $C\leq\mathbb{F}_{q}^{n\times m}$ be a nonzero rank-metric code with minimum rank distance $d$ and let $d^{\perp}$ be the rank distance of $C^{\perp}$. Then for every $U\leq\mathbb{F}_{q}^{n}$,
@@
There are some cases where we are able to say something more about the $q$-rank function of the $q$-polymatroid associated to a rank-metric code.
MRD codes constitute the most studied class of rank-metric codes. First of all, it is well known that MRD codes exist for every choice of parameters $n,m,d,q$. Furthermore, in [16] it is shown that if $m\geq n$, then the $q$-polymatroid arising from an MRD code is the uniform $q$-matroid of rank $n-d+1$, where $d$ is the minimum rank distance of the code. More specifically, we have the following result.
@@
[16, Cor. 6.6] Let $m\geq n$ and $C\leq\mathbb{F}_{q}^{n\times m}$ be an MRD code with minimum rank distance $d$. Let $\rho_{\mathcal{M}}$ be the rank function of the $q$-polymatroid $\mathcal{M}[C]$. Then for every $A\in\mathcal{L}(E)$,
@@
[12, Theorem 3.10] Let $m<n$ and $C\leq\mathbb{F}_{q}^{n\times m}$ be an MRD code with minimum rank distance $d$. Then for every $U\leq\mathbb{F}_{q}^{n}$, we have
@@
In particular, if $d>1$, $\mathcal{M}[C]$ is not a $q$-matroid; see [12, Cor. 3.11].

### v2

@@ 2. Preliminaries
The first results on $q$-polymatroids and their relation to coding theory can be read in [17] and [25]. The following definition is from [17, Def.4.1].
@@
The definition of $q$-polymatroid appeared in different forms in the literature. In [12, Def. 2.1] the same definition occurs with the only difference that the rank function takes only rational values. In this case we talk about rational $q$-polymatroids. Fundamental concepts regarding rational $q$-polymatroids have been investigated in [13] and [12]. A number $\mu\in\mathbb{Q}_{\geq 0}$ is called a denominator of $\rho$ and $\mathcal{M}$ if $\mu\rho(X)\in\mathbb{N}_{0}$ for all $X\in\mathcal{L}(E)$. The smallest denominator is called the principal denominator. A $q$-polymatroid with principal denominator $1$ (i.e., $\rho(X)\in\mathbb{N}_{0}$ for all $X$) is a $q$-matroid.
Next, for any $r\in\mathbb{N}$ a $(q,r)$-polymatroid as in [25, Def. 2] by Shiromoto, [10, Def. 1] and [3, Def. 1] can be turned into a rational $q$-polymatroid with denominator $r$ by dividing the rank function by $r$. Conversely, given a rational $q$-polymatroid with denominator $\mu$, then $(\mathcal{L}(E),\mu\rho)$ is a $(q,\lceil\mu\rceil)$-polymatroid. Finally, in [1, Def. 2.12] and [5, Def. 2.2] an $\mathcal{L}$-polymatroid is defined as a direct extension of the definition of the previous $q$-polymatroid.
We will use the following definition of independent spaces derived in [12, Def. 3.1].
@@
In [12, Def. 3.8], the authors define the $\mu$-bases of a space $V\in\mathcal{L}(E)$ as the inclusion-maximal $\mu$-independent subspaces of $V$. Furthermore, they show that all the $\mu$-bases of $V$ have the same rank value, which is equal to the rank of $V$.
@@
The following definition comes from [5, Def. 4.2].
@@ 4.2. The cyclic spaces
By following [15, Def. 7.7], we introduce the notion of full $q$-polymatroid.
@@ 4.3. The $\mu$-independent spaces
By definition of $\mathcal{M}$ we clearly have that the $\mu=\mathrm{denom}(\lambda)$ is always a denominator for $\mathcal{M}$, but in general it might not be principal. However, we have the following result that provides sufficient conditions for $\mu$ to be principal.
Theorem 4.12.
Let $\mu=\mathrm{denom}(\lambda)$ and assume that there exists a one-dimensional subspace $x\in\mathcal{L}(E)$ such that $x$ is strong independent in $\mathcal{M}$. Then $\mu$ is a principal denominator of $\mathcal{M}$.
@@
By [13, Rem. 2.7], we have that if $x\in\mathcal{L}(E)$ is strong independent in $\mathcal{M}$, then any denominator of $\mathcal{M}$ is integer. By the definition of $\mathcal{M}$ we clearly have that $\mu=\mathrm{denom}(\lambda)$ is always a denominator for $\mathcal{M}$. Since the independent spaces determine a $q$-matroid (see e.g. [4]), we see that there exists a space $X\in\mathcal{L}(E)$, such that $X$ is independent in $\mathcal{M}_{1}$, but not in $\mathcal{M}_{2}$. Then, there is a space $C\leq X$, which is a circuit in $\mathcal{M}_{2}$, but $C$ is independent in $\mathcal{M}_{1}$. Then $\rho_{2}(C)=\dim(C)-1$ and $\rho_{1}(C)=\dim(C)$. Let $\rho$ be the rank function of $\lambda\mathcal{M}_{1}+(1-\lambda)\mathcal{M}_{2}$. Then,
$\rho(C)=\lambda\rho_{1}(C)+(1-\lambda)\rho_{2}(C)=\rho_{2}(C)+\lambda(\rho_{1}(C)-\rho_{2}(C))=\rho_{2}(C)+\lambda.$
Then, $\mu=\mathrm{denom}(\lambda)$ is the smallest integer such that $\mu\rho(C)$ is an integer by the generalized euclidean lemma. Hence, we conclude that $\mu$ is the principal denominator of $\mathcal{M}$. ∎
@@ 5. The convex combination of paving q-matroids
[13, Prop 4.6.] Let $11\leq k\leq n-$ 1. In addition, let $\mathcal{S}\subseteq\mathcal{L}(E)_{k}$, such that $\dim(V\cap W)\leq k-2$, for every two distinct $V,W\in\mathcal{S}$. Define the map
@@
Therefore, if a $q$-polymatroid is paving according to Definition 5.4, then the dimension of every $\mu$-circuit is greater or equal than the rank of the $q$-polymatroid. A detailed discussion about $\mu$-independent spaces can be found in [12, Sec. 3-4].
@@ 7. The characteristic Puiseux polynomial
In [3, Sec. 3], the characteristic polynomial of an integer $(q,r)$-polymatroid was introduced and studied. Moreover the special case given by a $(q,1)$-polymatroid, i.e. a $q$-matroid, was then investigated in more detail in [20, Sec. 5]. In this section, we introduce a more general notion for rational $q$-polymatroids, inspired by [3]. However, this is not a polynomial anymore, but rather a truncated Puiseux series, i.e., a finite linear combination of monomials whose exponents may be rational and possibly negative. Afterwards, we discuss some properties in the context of $q$-polymatroids arising from convex combination of $q$-matroids.
@@
We want to emphasis that Theorem 7.4 does not say only that $\chi_{\mathcal{M}}$ is determined by either $\chi_{\mathcal{M}_{1}}$ or $\chi_{\mathcal{M}_{2}}$. By using [20, Thm. 5.8], we also observe that $\chi_{\mathcal{M}}$ can be computed by just using the flats of either $\mathcal{M}_{1}$ or $\mathcal{M}_{2}$. In other words, in order to determine $\chi_{\mathcal{M}}$ we do not need any concepts associated to the $q$-polymatroid $\mathcal{M}$, but rather only the well understood notion of flats of either of the $q$-matroids $\mathcal{M}_{1}$ or $\mathcal{M}_{2}$.
@@ 8. Representable q-polymatroid
Then $(\mathcal{L}(E),\rho)$ is a $q$-polymatroid [17, Theorem 5.3] and we denote it by $\mathcal{M}[C]$.
@@
Let $\rho:\mathcal{L}(E)\longrightarrow\mathbb{Z}_{\geq 0}$ be defined by $\rho(W):=k-\dim_{\mathbb{F}_{q^{m}}}(C(W)).$ Then $(\mathcal{L}(E),\rho)$ is a $q$-matroid [21, Thm. 24] and we also denote it by $\mathcal{M}[C]$. Note that in this case, $k$ is the dimension of $C$ over $\mathbb{F}_{q^{m}}$.
@@
[17, Prop. 6.2] Let $C\leq\mathbb{F}_{q}^{n\times m}$ be a nonzero rank-metric code with minimum rank distance $d$ and let $d^{\perp}$ be the rank distance of $C^{\perp}$. Then for every $U\leq\mathbb{F}_{q}^{n}$,
@@
There are some cases where we are able to say something more about the $q$-rank function of the $q$-polymatroid associated to a rank-metric code. MRD codes constitute the most studied class of rank-metric codes. First of all, it is well known that MRD codes exist for every choice of parameters $n,m,d,q$. Furthermore, in [17] it is shown that if $m\geq n$, then the $q$-polymatroid arising from an MRD code is the uniform $q$-matroid of rank $n-d+1$, where $d$ is the minimum rank distance of the code. More specifically, we have the following result.
@@
[17, Cor. 6.6] Let $m\geq n$ and $C\leq\mathbb{F}_{q}^{n\times m}$ be an MRD code with minimum rank distance $d$. Let $\rho_{\mathcal{M}}$ be the rank function of the $q$-polymatroid $\mathcal{M}[C]$. Then for every $A\in\mathcal{L}(E)$,
@@
[13, Theorem 3.10] Let $m<n$ and $C\leq\mathbb{F}_{q}^{n\times m}$ be an MRD code with minimum rank distance $d$. Then for every $U\leq\mathbb{F}_{q}^{n}$, we have
@@
In particular, if $d>1$, $\mathcal{M}[C]$ is not a $q$-matroid; see [13, Cor. 3.11].